determine the equation of the line passing through the point (-2,3) and perpendicular to the graph of x+y=24
Answer by venugopalramana(3286)

determine the equation of the line passing through the point (-2,3) and perpendicular to the graph of x+y=24
ANY PERPENDICULAR TO GIVEN LINE IS
X-Y = K..........IT PASSES THROUGH (-2,3)
HENCE
-2-3=K=-5
HENCE REQD.EQN.IS
X-Y=-5